### v1

# Bridging Conversational and Collaborative Signals for Conversational Recommendation

Abstract.
Conversational recommendation systems (CRS) leverage contextual information from conversations to generate recommendations but often struggle due to a lack of collaborative filtering (CF) signals, which capture user-item interaction patterns essential for accurate recommendations. We introduce Reddit-ML32M, a dataset that links reddit conversations with interactions on MovieLens 32M, to enrich item representations by leveraging collaborative knowledge and addressing interaction sparsity in conversational datasets. We propose an LLM-based framework that uses Reddit-ML32M to align LLM-generated recommendations with CF embeddings, refining rankings for better performance. We evaluate our framework against three sets of baselines: CF-based recommenders using only interactions from CRS tasks, traditional CRS models, and LLM-based methods relying on conversational context without item representations. Our approach achieves consistent improvements, including a 12.32% increase in Hit Rate and a 9.9% improvement in NDCG, outperforming the best-performing baseline that relies on conversational context but lacks collaborative item representations.
Conversational recommendation, large language models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Natural language generationccs: Computing methodologies Learning latent representations
1. Introduction
Figure 1. Illustration of the proposed CRS framework, which integrates LLM-based conversational recommendations with CF signals. The predictor uses these signals to generate ranked recommendations.
Conversational recommender systems (CRS) aim to elicit user preferences and generate recommendations through interactive dialogues (He et al., 2023; Li et al., 2018). Unlike traditional recommenders, CRS leverages conversational context to adapt to user preferences. However, existing CRS approaches face two main challenges: (1) They do not leverage collaborative filtering (CF) signals with conversational context, leading to suboptimal recommendations (Zhou et al., 2020; Friedman et al., 2023), and (2) CRS datasets are typically sparse, making it difficult to develop robust item representations for accurate recommendations. To this end, a typical CRS consists of two components: (1) a generator to produce natural language responses and (2) a recommender to rank items aligned with user preferences (He et al., 2023; Li et al., 2018; Chen et al., 2019).
From a CRS model perspective, CF-based approaches excel at modeling user-item relationships but fail to leverage conversational context (Salem et al., 2014; Kang and McAuley, 2018). Similarly, knowledge-driven methods improve semantic understanding using external knowledge graphs but lack the collaborative signals needed to model item relationships (Zhou et al., 2020). Large language models (LLMs) show promise in CRS due to their natural language understanding capabilities (He et al., 2023), but they rely at most on textual metadata, which constrains their ability to capture item relationships (He et al., 2025; Yang et al., 2024). Recent methods integrate CF embeddings with LLMs to improve recommendation tasks. For instance, A-LLMRec incorporates CF embeddings into prompts, while CoLLM maps them onto token embeddings (Kim et al., 2024; Zhang et al., 2023). ILM aligns CF embeddings with textual embeddings but lacks conversational context (Yang et al., 2024). While these methods incorporate CF with LLMs to improve recommendations, they are not designed to utilize conversational context. Our framework bridges this gap by integrating CF signals and conversational data into a unified CRS approach.
From CRS dataset perspective, current CRS datasets suffer from low interactions per item. For example, ReDIAL dataset (Li et al., 2018) contains 31,988 interactions, with 78.98% of items having fewer than 10 interactions (He et al., 2025). Similarly, Reddit-Movie (He et al., 2023) has 51,148 for 31,396 items as shown in Table 1. To our knowledge, no existing dataset bridges conversational context with CF information. To address these limitations, we propose a dataset that integrates CF signals with conversations and mitigates sparsity in CRS datasets.
In this work, we make two key contributions: First, we construct a new Reddit-ML32M dataset by linking the Reddit-Movie dataset (He et al., 2023) with the MovieLens 32M dataset (Harper and Konstan, 2015). This dataset enriches the interaction space, increasing the total number of interactions from 51,148 to over 30 million and improving dataset density from 0.013% to 0.68%, as shown in Table 1. The dataset facilitates the development of richer item representations for CRS tasks by exploiting both conversational data with CF information at the item level. Second, we propose a framework that integrates conversational context from LLMs with CF-based item representations as depicted in Figure 1. GPT-3.5t is used in a zero-shot setting as a backbone to generate initial recommendations from conversational context, which are then refined by ranking them against item embeddings generated by a pre-trained SASRec model (Kang and McAuley, 2018). This approach mitigates biases in LLM outputs, such as over-reliance on popular items or training data (He et al., 2023; Yang et al., 2024), while incorporating collaborative signals. Empirical evaluation across three paradigms: CF-based recommenders, CRS models, and LLM-based CRS without CF signals shows the effectiveness of our framework. Our approach achieves a 12.32% improvement in Hit Rate and a 9.9% improvement in NDCG over the best-performing baseline GPT-3.5t (He et al., 2025).
Table 1. Statistics for Reddit-Movie (He et al., 2023) and Reddit-ML32M datasets to improve item representations for CRS tasks.
Dataset #Interactions #Users #Items Density
Reddit 51,148 12,508 31,396 0.013%
Reddit-ML32M 30,074,259 200,947 22,014 0.68%
2. Method
Table 2. Comparison of performance against (1) CF-based recommenders, (2) CRS models, and (3) LLM-based CRS without CF signals for k = {1,5,10}. Results include percentage improvements over the best underlined values, with best values in bold.
Models H@1 N@1 H@5 N@5 H@10 N@10
PopRec .0022 ± .0004 .0022 ± .0004 .0087 ± .0011 .0046 ± .0004 .0131 ± .0011 .0064 ± .0005
item2vec .0078 ± .0009 .0078 ± .0009 .0274 ± .0016 .0131 ± .0006 .0436 ± .0020 .0172 ± .0017
FISM .0046 ± .0007 .0046 ± .0007 .0229 ± .0015 .0085 ± .0016 .0441 ± .0021 .0132 ± .0029
SASRec .0059 ± .0014 .0059 ± .0014 .0270 ± .0010 .0130 ± .0010 .0390 ± .0010 .0180 ± .0010
ReDIAL .0050 ± .0010 - .0290 ± .0010 .0190 ± .0010 .0440 ± .0010 .0240 ± .0010
UniCRS .0020 ± .0006 .0020 ± .0006 .0260 ± .0010 .0170 ± .0010 .0430 ± .0010 .0220 ± .0010
Llama2 - - .0420 ± .0010 .0270 ± .0010 .064 ± .0010 .0340 ± .0010
GPT-3.5-t .0152 ± .0012 .0152 ± .0012 .0686 ± .0024 .0289 ± .0012 .1218 ± .0031 .0419 ± .0013
Ours .0156 ± .0012 .0156 ± .0012 .0770 ± .0025 .0318 ± .0012 .1330 ± .0032 .0449 ± .0013
Improvement (%) +2.96% +2.96% +12.32% +9.91% +9.14% +7.22%
2.1. Problem Formulation
Let $\mathit{U}\coloneqq\{u_{1},\dots,u_{M}\}$ and $\mathit{I}\coloneqq\{v_{1},\dots,v_{N}\}$ represent sets of $M$ users and $N$ items, respectively. We define $\mathit{W}$ as the vocabulary of words forming an utterance $s\coloneqq(w_{i})_{i=1}^{m}$, with each $w_{i}\in\mathit{W}$.
Conversation Modality
We model a multi-turn conversation for recommendation tasks as $C\coloneqq(u^{t},s^{t},\mathit{I}^{t})_{t=1}^{T}$ with $T$ turns, where $\mathit{I}^{t}\subseteq\mathit{I}$ represents the items mentioned in turn $t$. For each turn, $s^{t}$ is the utterance, $u^{t}$ is the user in the conversation, and $\mathit{I}^{t}$ may be empty if no items are mentioned. A user seeking recommendations is a seeker, while a recommender provides recommendations.
Interaction Modality
Observed user-item interactions are defined as $\mathit{R}\coloneqq\{(u,v)\mid u\in\mathit{U},v\in\mathit{I},\text{ and }\mathbf{1}(u,v)=1\}$, where $\mathbf{1}(u,v)=1$ if user $u$ interacted with item $v$. For conversations, $\mathit{R}$ is:
(1) $\mathit{R}=\{(u,v)\mid u\in\mathit{U},\exists t\text{ such that }v\in\mathit{I}^{t}\text{ and }\mathbf{1}(u,v,t)=1\}$
where $\mathbf{1}(u,v,t)$ indicates if seeker $u$ mentions item $v$ at turn $t$. However, CRS datasets often lack sufficient user-item interactions for learning robust collaborative signals, as shown in Table 1  [^1] Density is calculated as #Interactions/(#Users + #Items).
Linking with Historical Interaction Dataset
To address the user-item interactions limitations of conversational datasets, we create a linked historical interaction dataset $\mathit{D}_{CF}=(\mathit{X},\mathit{V},\mathit{Z})$ to provide collaborative signals for learning item representations. The dataset consists of three components: (1) $\mathit{X}$ is the set of users in $\mathit{D}_{CF}$, and (2) $\mathit{V}$ is the set of items in $\mathit{D}_{CF}$, with $\mathit{V}\subseteq\mathit{I}$ to ensure alignment with the items in the conversational dataset, and (3) $\mathit{Z}$ is the set of user-item interaction sequences. Each user $x\in\mathit{X}$ has an interaction sequence $Z^{x}=(i^{x}_{1},i^{x}_{2},\ldots,i^{x}_{k})\in\mathit{Z}$, where $i^{x}_{k}\in\mathit{V}$ denotes the $k$-th item interacted with by user $x$. These sequences serve as a foundation for learning collaborative signals to generate rich item representations transferable to CRS.
Sequential Recommendation for Item Representation Learning
The goal of sequential recommendation is to predict the next item in a user interaction sequence based on their historical interactions. Given the interaction sequences $\mathit{Z}$, for a user $x\in\mathit{X}$, let $Z^{x}_{1:k}=(i^{x}_{1},i^{x}_{2},\ldots,i^{x}_{k})$ be the first $k$ items of the sequence. The item embeddings are represented by the matrix $E\in\mathbb{R}^{|V|\times d}$ as:
(2) $E^{x}_{1:k}=(E_{i^{x}_{1}},E_{i^{x}_{2}},\ldots,E_{i^{x}_{k}})\in\mathbb{R}^{k\times d}$
This sequence embedding matrix is input into a sequential recommender to predict a next item $i^{x}_{k+1}$. The training objective is maximizing likelihood of predicting the next item:
(3) $\max_{\Theta}\sum_{x\in\mathit{X}}\sum_{k=1}^{|Z^{x}|-1}\log p(i^{x}_{k+1}\mid Z^{x}_{1:k};\Theta)$
where $p(i^{x}_{k+1}\mid Z^{x}_{1:k};\Theta)$ represents the probability of the $(k+1)$-th interaction, conditioned on the previous $k$ items, and $\Theta$ is the set of model parameters. By optimizing Equation 3, the model generates item embeddings $E$ that encode collaborative signals. Since $\mathit{V}\subseteq\mathit{I}$, embeddings can be directly applied for the conversational dataset to enable effective item ranking and recommendations. Note that, while our framework focuses on sequential models for generating item embeddings, it is adaptable to non-sequential settings by replacing the backbone model with non-sequential alternatives such as FISM (Kabbur et al., 2013). This flexibility allows the framework to support a range of recommendation paradigms.
Objective
The recommender component, following representative works (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020; He et al., 2023), optimizes item selection during each conversation turn. During the $k$-th turn, the component exploits the preceding context $(u^{t},s^{t},\mathit{I}^{t})_{t=1}^{k-1}$ to generate a ranked list $\hat{\mathit{I}}^{k}$, which aligns with ground-truth items in $\mathit{I}^{k}$.
2.2. Proposed Framework
We propose a framework to improve CRS as illustrated in Figure 1. Unlike previous models focusing on conversational contexts (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020; Wang et al., 2022; He et al., 2023; Deng et al., 2021), we introduce dataset links between CRS and CF modalities to integrate user interaction information from both textual content and item representations generated via implicit feedback. Additionally, we incorporate LLMs in a zero-shot setting (He et al., 2023; Friedman et al., 2023) leveraging item representations generated by a pretrained CF recommender as supplementary information.
Prompting
For each conversation, the context $(u^{t},s^{t},\mathit{I}^{t})_{t=1}^{k-1}$ is used to prompt the LLM, with the $k$-th turn being the recommender’s reply. $s^{t}$ contains raw text for processing by the LLM. A prompt specifies that the input is intended for conversational recommendation. Pre-trained LLMs are employed in a zero-shot setting to generate item predictions (Radford et al., 2019; He et al., 2023). prompt structures the input as:
[…]
where prompt is the constructed prompt, [TASK] is a task description template, [FORMAT] is format requirements, and [CONTEXT] is the conversational context. Figure 1 illustrates the transformation.
LLMs
We employ LLMs in a zero-shot setting to generate recommendations from prompt (He et al., 2023). For reproducibility, we configure deterministic inference by setting the temperature parameter $\tau=0$, ensuring the softmax selects the highest-probability word at each step (Ackley et al., 1985). The response $=(w_{1},w_{2},\dots,w_{n})$ generated as:
$w_{i}=\arg\max_{j}\phi_{j}\quad\text{for }i=1,2,\dots,n$
where $\phi_{j}$ is the logit for candidate word $j$ at step $i$. This approach leverages the pre-trained knowledge of LLMs without fine-tuning.
Post-LLM Item Retrieval
Rather than evaluating model weights or output logits from LLMs directly, we implement an exact matching post-processing step that maps a list of recommendations in natural language into a set of in-dataset items $\mathit{I}_{LLM}$.
Embedding-Based Item Ranking
To refine the LLM-generated recommendations $\mathit{I}_{LLM}$, we compute a pairwise similarity matrix between the embeddings of $\mathit{I}_{LLM}$ and all items in the dataset $\mathit{V}$. Using the pretrained embedding matrix $E\in\mathbb{R}^{|V|\times d}$ from SASRec, where $d$ is the embedding dimensionality, we first retrieve the embeddings of the LLM-recommended items, $E_{LLM}\in\mathbb{R}^{m\times d}$. The pairwise similarity matrix $S\in\mathbb{R}^{m\times|V|}$ is computed as:
(5) $S_{i,j}=\frac{\mathbf{e}_{i}\cdot\mathbf{e}_{j}}{\|\mathbf{e}_{i}\|\|\mathbf{e}_{j}\|},\quad\forall i\in\mathit{I}_{LLM},\,j\in\mathit{V},$
where $\mathbf{e}_{i}$ and $\mathbf{e}_{j}$ are the embeddings of items $i$ and $j$, respectively. The matrix contains similarity scores between each LLM recommended item and all items in $\mathit{V}$. We use max pooling across the rows of $S$ to aggregate scores for each item in $\mathit{V}$:
(6) $s(j)=\max_{i\in\mathit{I}_{LLM}}S_{i,j},\quad\forall j\in\mathit{V}.$
This results in a score vector $\mathbf{s}\in\mathbb{R}^{|V|}$, where $s(j)$ represents the maximum similarity of item $j\in\mathit{V}$ with any item in $\mathit{I}_{LLM}$. Items in $\mathit{V}$ are ranked based on their scores as $\hat{\mathit{I}}^{k}=\text{argsort}_{j\in\mathit{V}}(-s(j))$.
3. Experiments
Experimental Setup
Our framework integrates embeddings generated by SASRec (Kang and McAuley, 2018) for collaborative filtering with conversational recommendations generated by GPT-3.5-t to create a unified CRS that leverages the strengths of both approaches. To assess performance, we evaluate our framework against three groups of representative baseline models. First, traditional item-based recommendation models include popularity (PopRec), FISM (Kabbur et al., 2013), item2vec (Barkan and Koenigstein, 2016), and SASRec (Kang and McAuley, 2018), trained on interactions from the Reddit Movie dataset, as shown in Table 1. These models serve as representative baselines for assessing recommendation performance. Second, traditional CRS models include ReDIAL [^2] Results as reported in (He et al., 2023, 2025)(Li et al., 2018) and UniCRS (Wang et al., 2022). UniCRS leverages pre-trained language models with prompt tuning to generate recommendations. We do not other works (Li et al., 2022; Ma et al., 2021), as UniCRS is representative with similar results. Third, zero-shot LLM-based CRS models, including LLaMA2-7B [^2] Results as reported in (He et al., 2023, 2025) (Touvron et al., 2023) and GPT-3.5-t, which generate conversational recommendations without fine-tuning (He et al., 2023). GPT-3.5-t stands out as a state-of-the-art model for CRS tasks, as noted in (He et al., 2025) under the zero-shot setting defined in (He et al., 2023). Our experiments confirm that GPT-3.5-t achieves the best results across key metrics, validating its position as a leading baseline for CRS.
Evaluation Protocol
Following prior studies (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020; He et al., 2023; Wang et al., 2022; He et al., 2025), we generate predictions for each conversation and select the top $k$ recommendations. We evaluate against ground truth items $\hat{\mathit{I}}^{k}$. We report Hit Rate (H@$k$) and NDCG@$k$ (N@$k$) for $K=\{1,5,10\}$.
3.1. Results and Analysis
We assess performance of our framework against eight baselines as shown in Table 2. Our method consistently outperforms all baselines, demonstrating the effectiveness of integrating CF signals into conversational recommendations. Specifically, our framework achieves a 12.32% improvement in H@5 over the state-of-the-art baseline (He et al., 2025), GPT-3.5-t. Similarly, N@5 improves by 9.91%, further confirming the ranking quality of our method.
CF-based models perform well in leveraging historical user-item interactions but fall short in conversational settings due to their inability to incorporate contextual information. In contrast, CRS models such as ReDIAL and UniCRS incorporate conversational data but are limited in their ability to capture behavioral patterns. LLM-based CRS models leverage conversational context to achieve strong performance but tend to fail to generalize to less frequent items due to the absence of item representations. Our framework addresses these limitations by using CF embeddings with LLM-generated recommendations, resulting in superior accuracy across all scenarios. Unlike SASRec, which relies on historical interactions, or GPT-3.5-t, which lacks item-level embeddings, our framework bridges the gap between CF and conversational understanding.
4. Related Work
Exploiting collaborative signals can be done using matrix factorization (Koren et al., 2009) or a Graph Convolutional Network (GCN) (He et al., 2020). SASRec (Kang and McAuley, 2018) uses a self-attention network to exploit sequential patterns within collaborative signals. Using conversational context commonly involves exploiting the semantics of conversational text and recommending items in order to maximize alignment between conversation and item attributes. ReDial (Li et al., 2018) predicts user sentiment in the conversational context and feeds it to an autoencoder-based recommendation. EAR (Lei et al., 2020) extracts item attributes and attribute-specific user preferences from the conversational context and feeds them to an attribute-aware Bayesian Personalized Ranking (BPR) model for recommendation. KBRD (Chen et al., 2019) extracts item and non-item entities from the conversation and feeds them through a Relational Graph Convolutional Network (RGCN) to generate a knowledge-enhanced user preference representation. KGSF (Zhou et al., 2020) exploits a word-oriented knowledge graph via a GCN in addition to item-oriented knowledge to recommend items based on keywords in the conversation. UniCRS (Wang et al., 2022) jointly optimizes entity embeddings from an RGCN and contextual word embeddings from pretrained language models. Recent works have explored the direct use of conversational context by prompting LLMs (He et al., 2023; Friedman et al., 2023). ILM (Yang et al., 2024) uses item-text contrastive learning for learning text-aligned item representations from collaborative signals. (Zhang et al., 2023) proposes tuning a low-rank adapter associated with LLM and a collaborative information encoding module so that collaborative information can be exploited by LLMs. TALLRec (Bao et al., 2023) proposes instruction tuning of LLMs to improve their ability to exploit collaborative signals.
5. Conclusion and Future Directions
We present the Reddit-ML32M dataset and a framework that integrates CF signals with conversational context to improve CRS. Our approach achieves consistent performance improvements, with a 12.32% increase in recall and a 9.91% improvement in NDCG compared to existing methods. The dataset serves as a resource for research in CRS, while our methodology shows promising future directions for constructing diverse datasets across different domains. This opens opportunities for broader CRS applications and tackling interaction sparsity in future work.

### v2

@@
 Conversational recommendation, large language models
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Natural language generationccs: Computing methodologies Learning latent representations
+journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715486isbn: 979-8-4007-1331-6/25/04ccs: Information systems Recommender systemsccs: Computing methodologies Natural language generationccs: Computing methodologies Learning latent representations
 1. Introduction
 Figure 1. Illustration of the proposed CRS framework, which integrates LLM-based conversational recommendations with CF signals. The predictor uses these signals to generate ranked recommendations.
-Conversational recommender systems (CRS) aim to elicit user preferences and generate recommendations through interactive dialogues (He et al., 2023; Li et al., 2018).
+Conversational recommender systems (CRS) aim to elicit user preferences and generate recommendations through interactive dialogues (He et al., 2023; Li et al., 2018; Sun and Zhang, 2018; Wang et al., 2023).
 Unlike traditional recommenders, CRS leverages conversational context to adapt to user preferences.
@@
 Similarly, knowledge-driven methods improve semantic understanding using external knowledge graphs but lack the collaborative signals needed to model item relationships (Zhou et al., 2020).
-Large language models (LLMs) show promise in CRS due to their natural language understanding capabilities (He et al., 2023), but they rely at most on textual metadata, which constrains their ability to capture item relationships (He et al., 2025; Yang et al., 2024).
+Large language models (LLMs) show promise in CRS due to their natural language understanding capabilities (Brown et al., 2020; Chowdhery et al., 2023; Ouyang et al., 2022), but they rely at most on textual metadata, which constrains their ability to capture item relationships (He et al., 2025; Yang et al., 2024).
 Recent methods integrate CF embeddings with LLMs to improve recommendation tasks.
-For instance, A-LLMRec incorporates CF embeddings into prompts, while CoLLM maps them onto token embeddings (Kim et al., 2024; Zhang et al., 2023).
+For instance, A-LLMRec incorporates CF embeddings into prompts, while CoLLM maps them onto token embeddings (Kim et al., 2024; Zhang et al., 2023a).
 ILM aligns CF embeddings with textual embeddings but lacks conversational context (Yang et al., 2024).
@@
 Second, we propose a framework that integrates conversational context from LLMs with CF-based item representations as depicted in Figure 1.
-GPT-3.5t is used in a zero-shot setting as a backbone to generate initial recommendations from conversational context, which are then refined by ranking them against item embeddings generated by a pre-trained SASRec model (Kang and McAuley, 2018).
+GPT-3.5-turbo (OpenAI, 2024) [^1] We refer to this model as GPT-3.5t hereafter is used in a zero-shot setting as a backbone to generate initial recommendations from conversational context, which are then refined by ranking them against item embeddings generated by a pre-trained SASRec model (Kang and McAuley, 2018).
 This approach mitigates biases in LLM outputs, such as over-reliance on popular items or training data (He et al., 2023; Yang et al., 2024), while incorporating collaborative signals.
@@
 2. Method
-Table 2.
-Comparison of performance against (1) CF-based recommenders, (2) CRS models, and (3) LLM-based CRS without CF signals for k = {1,5,10}.
-Results include percentage improvements over the best underlined values, with best values in bold.
-Models H@1 N@1 H@5 N@5 H@10 N@10
-PopRec .0022 ± .0004 .0022 ± .0004 .0087 ± .0011 .0046 ± .0004 .0131 ± .0011 .0064 ± .0005
-item2vec .0078 ± .0009 .0078 ± .0009 .0274 ± .0016 .0131 ± .0006 .0436 ± .0020 .0172 ± .0017
-FISM .0046 ± .0007 .0046 ± .0007 .0229 ± .0015 .0085 ± .0016 .0441 ± .0021 .0132 ± .0029
-SASRec .0059 ± .0014 .0059 ± .0014 .0270 ± .0010 .0130 ± .0010 .0390 ± .0010 .0180 ± .0010
-ReDIAL .0050 ± .0010 - .0290 ± .0010 .0190 ± .0010 .0440 ± .0010 .0240 ± .0010
-UniCRS .0020 ± .0006 .0020 ± .0006 .0260 ± .0010 .0170 ± .0010 .0430 ± .0010 .0220 ± .0010
-Llama2 - - .0420 ± .0010 .0270 ± .0010 .064 ± .0010 .0340 ± .0010
-GPT-3.5-t .0152 ± .0012 .0152 ± .0012 .0686 ± .0024 .0289 ± .0012 .1218 ± .0031 .0419 ± .0013
-Ours .0156 ± .0012 .0156 ± .0012 .0770 ± .0025 .0318 ± .0012 .1330 ± .0032 .0449 ± .0013
-Improvement (%) +2.96% +2.96% +12.32% +9.91% +9.14% +7.22%
 2.1. Problem Formulation
@@
 During the $k$-th turn, the component exploits the preceding context $(u^{t},s^{t},\mathit{I}^{t})_{t=1}^{k-1}$ to generate a ranked list $\hat{\mathit{I}}^{k}$, which aligns with ground-truth items in $\mathit{I}^{k}$.
+2.2. Dataset
+We introduce Reddit-ML32M, a novel dataset that bridges Reddit Movie dataset for CRS tasks (He et al., 2023) with MovieLens 32M, a collaborative filtering dataset (Harper and Konstan, 2015).
+The key challenge is aligning items across these datasets.
+To address this, we use a two-phase linking method.
+First, IMDb IDs serve as unique identifiers, as they exist in both datasets.
+However, direct matching may fail when IMDb IDs are modified, reassigned, or deprecated over time, leading to inconsistencies between datasets.
+To resolve these cases, we use IMDbPY  [^3] https://github.com/MaximShidlovski23/imdbpy to retrieve historical IMDb IDs using exact title and release date matching, ensuring correct alignment even when an item’s IMDb ID has changed over time or been deprecated.
+This linking process enriches the interaction space, increasing total interactions from 51,148 to over 30 million, with a corresponding density improvement from 0.013% to 0.68% as shown in Table 1.
+We partition the dataset into 80% training, 10% validation, and 10% test sets.
 2.3. Proposed Framework
-We propose a framework to improve CRS as illustrated in Figure 1.
+We propose BridgeCRS, a framework designed to improve CRS, as illustrated in Figure 1 and outlined in Algorithm 1.
 Unlike previous models focusing on conversational contexts (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020; Wang et al., 2022; He et al., 2023; Deng et al., 2021), we introduce dataset links between CRS and CF modalities to integrate user interaction information from both textual content and item representations generated via implicit feedback.
@@
 Items in $\mathit{V}$ are ranked based on their scores as $\hat{\mathit{I}}^{k}=\text{argsort}_{j\in\mathit{V}}(-s(j))$.
+Algorithm 1 BridgeCRS Algorithm
+1:Conversation history $C$; pre-trained LLM $\mathcal{M}$; CF embedding matrix $E$; Candidate item set $\mathit{V}$
+2:Ranked recommendation list $\hat{\mathit{I}}^{k}$
+3:LLM-Based Recommendation
+4:$\texttt{response}\leftarrow\mathcal{M}(\mathit{F}(\texttt{[TASK]},\texttt{[FORMAT]},C))$
+5:Extract $\mathit{I}_{LLM}$ from response
+6:Post-LLM Item Retrieval
+7:$\mathit{I}_{LLM}\leftarrow\mathit{I}_{LLM}\cap\mathit{V}$triangleright Keep only items in candidate set
+8:Embedding-Based Item Ranking
+9:Retrieve embeddings $E_{LLM}\in\mathbb{R}^{m\times d}$ for $\mathit{I}_{LLM}$
+10:Compute $s(j)=\max_{i\in\mathit{I}_{LLM}}\frac{\mathbf{e}_{i}\cdot\mathbf{e}_{j}}{\|\mathbf{e}_{i}\|\|\mathbf{e}_{j}\|},\quad\forall j\in\mathit{V}$
+11:return $\hat{\mathit{I}}^{k}=\text{argsort}_{j\in\mathit{V}}(-s(j))$
 3. Experiments
+Table 2.
+Comparison of performance against (1) CF-based recommenders, (2) CRS models, and (3) LLM-based CRS without CF signals for k = {1,5,10}.
+Results include percentage improvements over the best underlined values, with best values in bold.
+Models H@1 N@1 H@5 N@5 H@10 N@10
+PopRec .0022 ± .0004 .0022 ± .0004 .0087 ± .0011 .0046 ± .0004 .0131 ± .0011 .0064 ± .0005
+item2vec .0078 ± .0009 .0078 ± .0009 .0274 ± .0016 .0131 ± .0006 .0436 ± .0020 .0172 ± .0017
+FISM .0046 ± .0007 .0046 ± .0007 .0229 ± .0015 .0085 ± .0016 .0441 ± .0021 .0132 ± .0029
+SASRec .0059 ± .0014 .0059 ± .0014 .0270 ± .0010 .0130 ± .0010 .0390 ± .0010 .0180 ± .0010
+ReDIAL .0050 ± .0010 - .0290 ± .0010 .0190 ± .0010 .0440 ± .0010 .0240 ± .0010
+UniCRS .0020 ± .0006 .0020 ± .0006 .0260 ± .0010 .0170 ± .0010 .0430 ± .0010 .0220 ± .0010
+Llama2 - - .0420 ± .0010 .0270 ± .0010 .064 ± .0010 .0340 ± .0010
+GPT-3.5-t .0152 ± .0012 .0152 ± .0012 .0686 ± .0024 .0289 ± .0012 .1218 ± .0031 .0419 ± .0013
+BridgeCRS (Ours) .0156 ± .0012 .0156 ± .0012 .0770 ± .0025 .0318 ± .0012 .1330 ± .0032 .0449 ± .0013
+Improvement (%) +2.96% +2.96% +12.32% +9.91% +9.14% +7.22%
 Experimental Setup
@@
 Similarly, N@5 improves by 9.91%, further confirming the ranking quality of our method.
+Note that H@1 and N@1 are identical because NDCG@1 reduces to Hit@1 due to the absence of rank discounting.
 CF-based models perform well in leveraging historical user-item interactions but fall short in conversational settings due to their inability to incorporate contextual information.
 In contrast, CRS models such as ReDIAL and UniCRS incorporate conversational data but are limited in their ability to capture behavioral patterns.
-LLM-based CRS models leverage conversational context to achieve strong performance but tend to fail to generalize to less frequent items due to the absence of item representations.
+LLM-based CRS models leverage conversational context effectively but struggle with less frequent items due to the absence of item representations.
 Our framework addresses these limitations by using CF embeddings with LLM-generated recommendations, resulting in superior accuracy across all scenarios.
 Unlike SASRec, which relies on historical interactions, or GPT-3.5-t, which lacks item-level embeddings, our framework bridges the gap between CF and conversational understanding.
+Generalization to Other Domains
+The proposed dataset linking method is broadly applicable across domains where structured identifiers facilitate dataset alignment, such as ASINs in e-commerce and ISBNs in books (Ni et al., 2019; Wan and McAuley, 2018).
+This approach extends to industry settings where proprietary user interactions contain unique product identifiers, allowing collaborative item representations to be generated from implicit feedback sources such as purchase history and browsing patterns (Yoon et al., 2024).
+By leveraging structured identifiers and user interaction data, this framework can integrate collaborative signals into CRS tasks, which bridging structured and conversational modalities across diverse domains.
 4. Related Work
-Exploiting collaborative signals can be done using matrix factorization (Koren et al., 2009) or a Graph Convolutional Network (GCN) (He et al., 2020).
+Exploiting collaborative signals can be achieved using matrix factorization (Koren et al., 2009) or a Graph Convolutional Network (GCN) (He et al., 2020).
 SASRec (Kang and McAuley, 2018) uses a self-attention network to exploit sequential patterns within collaborative signals.
-Using conversational context commonly involves exploiting the semantics of conversational text and recommending items in order to maximize alignment between conversation and item attributes.
+Using conversational context typically involves leveraging the semantics of conversational text and recommending items to maximize alignment between conversation and item attributes.
 ReDial (Li et al., 2018) predicts user sentiment in the conversational context and feeds it to an autoencoder-based recommendation.
-EAR (Lei et al., 2020) extracts item attributes and attribute-specific user preferences from the conversational context and feeds them to an attribute-aware Bayesian Personalized Ranking (BPR) model for recommendation.
-KBRD (Chen et al., 2019) extracts item and non-item entities from the conversation and feeds them through a Relational Graph Convolutional Network (RGCN) to generate a knowledge-enhanced user preference representation.
+The autoencoder-based recommender is similar to AutoRec (Sedhain et al., 2015) which is trained to produce unobserved user ratings from the observed user ratings.
+EAR (Lei et al., 2020) extracts item attributes and attribute-specific user preferences from the dialogue context and feeds them to an attribute-aware Bayesian Personalized Ranking (BPR) model for recommendation.
+KBRD (Chen et al., 2019) extracts item and non-item entities from the conversation and feeds them through a Relational Graph Convolutional Network (RGCN) (Schlichtkrull et al., 2018) to generate a knowledge-enhanced user preference representation.
+The RGCN is trained with an item-oriented knowledge graph from DBpedia (Lehmann et al., 2015).
 KGSF (Zhou et al., 2020) exploits a word-oriented knowledge graph via a GCN in addition to item-oriented knowledge to recommend items based on keywords in the conversation.
+The GCN associated with the word-oriented knowledge graph is trained on ConceptNet (Speer et al., 2017).
 UniCRS (Wang et al., 2022) jointly optimizes entity embeddings from an RGCN and contextual word embeddings from pretrained language models.
-Recent works have explored the direct use of conversational context by prompting LLMs (He et al., 2023; Friedman et al., 2023).
+Recent works have explored the direct use of conversational context by prompting LLMs (Friedman et al., 2023; Hou et al., 2024).
+He et al.
+(He et al., 2023) feed dialogue context directly to LLMs for generating recommendations and do not use user-item interaction data.
+Liu et al.
+(Liu et al., 2023) demonstrate the use of LLMs for several recommendation tasks such as direct recommendation, rating prediction, sequential recommendation, explanation generation, and review summarization by constructing task-specific prompts.
+However, this approach performs poorly on direct and sequential recommendation (Zhang et al., 2023b).
+Gao et al.
+(Gao et al., 2023) propose adding user-item interaction history to the LLM prompts in addition to the dialogue context.
+However, LLMs cannot exploit complex user-item interactions from their textual representations.
 ILM (Yang et al., 2024) uses item-text contrastive learning for learning text-aligned item representations from collaborative signals.
-(Zhang et al., 2023) proposes tuning a low-rank adapter associated with LLM and a collaborative information encoding module so that collaborative information can be exploited by LLMs.
+They train an item encoder to produce text-aligned item representations so that a frozen LLM may exploit item representations with existing parametric knowledge.
+Zhang et al.
+(Zhang et al., 2023a) propose encoding user-item interaction using traditional collaborative models and tuning a low-rank adapter associated with LLM so that collaborative information can be exploited by LLMs.
 TALLRec (Bao et al., 2023) proposes instruction tuning of LLMs to improve their ability to exploit collaborative signals.